# A Robust and Non-Iterative Tensor Decomposition Method with Automatic Thresholding

### v2

@@ 5 Discussion
TARST consistently achieved the lowest RRSE across all experimental settings and maintained stable performance even under increasing noise levels and outlier ratios. This robustness stems from its automatic thresholding mechanism based on the Marčenko–Pastur (MP) distribution, which dynamically adjusts the threshold according to the noise intensity, removing noise-dominated singular values while preserving true signal components. Consequently, the error growth remains nearly linear with respect to $\sigma$, and higher-dimensional settings further enhance stability due to the concentration of the singular-value distribution around its expectation. Because TARST automatically determines the statistically valid rank for each mode, it avoids the under- and over-estimation errors frequently observed in fixed-rank methods such as HOSVD and HOOI. Moreover, its non-iterative structure eliminates sensitivity to initialization and local minima, achieving both computational efficiency and stability. Overall, by combining automatic rank estimation with a non-iterative design, TARST mitigates the bias of fixed-rank methods and the variance of iterative ones, maintaining consistent robustness against noise and scale variations.
@@
In the outlier experiments (Figure 3), RRSE surfaces reveal that HOSVD and HOOI exhibit substantial error increases as the outlier ratio or outlier scale increases. In contrast, Tucker-L2E demonstrates partial robustness, maintaining low RRSE under moderate contamination and achieving lower error than other methods around $r=20\%$, where its loss function mitigates the influence of moderate outliers. However, as contamination increases (e.g., $r>25\%$), its performance degrades. TARST, by comparison, maintains uniformly low RRSE values across nearly all conditions, showing minimal variation even under extreme settings (e.g., $r=50\%$, scale = 100). This indicates that the proposed thresholding strategy limits the influence of extreme observations by discarding outlier-dominated singular components, resulting in stable reconstruction performance across contamination conditions. These results suggest that TARST can adapt to situations where a small fraction of data exhibits highly dispersed values—analogous to heavy-tailed noise frequently encountered in real-world datasets such as financial transactions, energy demand, or sensor readings. Unlike fixed-rank or iterative approaches, TARST automatically adjusts its rank to exclude such fluctuations without an explicit noise model. This behavior implies that the MP-based thresholding functions as a robustness regularizer, suppressing extreme singular components while preserving the global multilinear structure. Consequently, TARST demonstrates both noise resistance and resilience to heavy-tailed perturbations, bridging statistical robustness and computational tractability in tensor decomposition.
From a theoretical standpoint, TARST generalizes the optimal thresholding principle of the MP law to tensor data, providing the first statistically grounded framework for automatic rank determination in tensor decomposition. Inspired by Absi and Heuvel, (2019), this study also applied WCTR analysis to tensor computation, demonstrating that TARST’s computational complexity is bounded by a polynomial order $\mathcal{O}(P^{1+\varepsilon})$ with $\varepsilon<1$. This result guarantees tractable computation for high-dimensional tensor analysis and introduces a new analytical framework for evaluating algorithmic efficiency in Operations Research.
@@
From a practical perspective, TARST’s non-iterative and parameter-free design ensures reproducible results independent of analyst expertise or initialization. Its robustness to noise and outliers makes it applicable to various OR problems such as demand forecasting, anomaly detection, and efficiency evaluation involving multidimensional data. Furthermore, its simplicity and scalability enable deployment in large-scale and real-time analytical environments, including IoT and smart city monitoring.
@@
A current limitation of TARST is the assumption of i.i.d. Gaussian noise. Future research should extend the framework to handle non-Gaussian, anisotropic, or time-varying noise conditions. Algorithmically, developing distributed or online implementations and incorporating adaptive mode-wise thresholding would enhance scalability for streaming and large-scale data. From an application viewpoint, applying TARST to real-world datasets in domains such as transportation, finance, and energy, and integrating it with optimization-based decision-support systems, represent promising directions. Additionally, establishing asymptotic guarantees for the estimated thresholds and comparing TARST’s statistical efficiency with other robust estimators would further strengthen its theoretical foundation. In summary, TARST achieves both theoretical optimality and computational efficiency while overcoming key limitations of conventional tensor decomposition methods, including rank specification, iterative optimization, and initialization sensitivity. The proposed framework offers a robust, reproducible, and theoretically grounded foundation for high-dimensional data analysis, contributing to the advancement of data-driven decision support in Operations Research.
@@ 6 Conclusion
This study proposed TARST, a non-iterative and rank-free tensor decomposition method, to overcome two major limitations of conventional approaches: the need for pre-specified ranks and the computational instability caused by iterative optimization. By employing automatic thresholding based on the MP distribution, TARST achieves statistically valid rank estimation. Through theoretical analysis using the WCTR, the method was shown to guarantee computational tractability within polynomial time. Accordingly, TARST provides a new methodological foundation for tensor decomposition that achieves both theoretical optimality and computational efficiency.
Simulation experiments demonstrated that TARST maintained the lowest reconstruction error under various noise intensities and outlier ratios, showing remarkable stability compared with existing methods. These results indicate that the proposed approach is consistently robust to noise, outliers, and scale variations. In particular, its ability to estimate the underlying structure without rank specification or iterative refinement offers significant practical advantages.
From the perspective of Operations Research, TARST contributes to both methodological development and practical decision support. Methodologically, it establishes a statistically grounded and computationally tractable framework for high-dimensional tensor analysis, bridging the gap between statistical estimation theory and algorithmic efficiency in OR. Practically, the robustness and parameter-free nature of TARST enable reliable, analyst-independent analysis in data-driven decision-support contexts, such as demand forecasting, anomaly detection, and performance evaluation. Thus, TARST serves as a foundational analytical tool that connects theoretical rigor with operational applicability.
Future research directions include extending the method to handle non-Gaussian or time-varying noise, developing distributed and online implementations, and integrating TARST with optimization algorithms to realize real-time decision-support systems. Through these extensions, TARST has the potential to further bridge theory and practice in high-dimensional data analysis and to contribute to the continued advancement of Operations Research.

### v3

@@ 5 Discussion
The theoretical analysis confirmed that TARST achieves tractable computational scalability, operating within polynomial time $\mathcal{O}(P^{1+\varepsilon})$ under two general conditions: (i) the post-threshold ranks of each mode grow more slowly than their corresponding mode dimensions ($\alpha_{k}<1$), and (ii) the tensor order $N$ is finite. Under these assumptions, the exponent $\varepsilon=\max{1/2,\alpha_{\max}}$ remains strictly below one, indicating that the computational complexity exponent is smaller than two. This result theoretically guarantees stable scalability even for high-order or anisotropic tensor data and explains why the proposed method maintains tractable computation as the tensor dimension or data volume increases. In addition to computational efficiency, TARST generalizes the optimal singular-value thresholding principle of the MP law to the tensor domain, providing the first statistically grounded framework for automatic rank determination in tensor decomposition. Moreover, by incorporating the WCTR analysis framework Absi and Heuvel, (2019), this study introduces a new theoretical perspective for assessing algorithmic efficiency in tensor computation, bridging statistical optimality with computational tractability in OR.
@@
Experimentally, TARST consistently achieved the lowest RRSE across all settings and maintained stable performance under increasing noise levels and outlier ratios. This robustness arises from its automatic thresholding mechanism based on the MP distribution, which dynamically adjusts the threshold according to noise intensity, removing noise-dominated singular values while preserving true signal components. Consequently, error growth remains nearly linear with respect to $\sigma$, and higher-dimensional settings further enhance stability due to the concentration of the singular-value distribution around its expectation. Because TARST automatically determines the statistically valid rank for each mode, it avoids the under- and overestimation errors frequently observed in fixed-rank methods such as HOSVD and HOOI. Moreover, its non-iterative structure removes dependence on initialization and local minima, ensuring both computational efficiency and stability. Overall, by combining automatic rank estimation with a non-iterative design, TARST mitigates the bias of fixed-rank methods and the variance of iterative ones, maintaining consistent robustness against noise and scale variation.
@@
In the outlier experiments (Figure 3), RRSE surfaces reveal that HOSVD and HOOI exhibit substantial error increases as the outlier ratio or outlier scale increases. In contrast, Tucker-L2E demonstrates partial robustness, maintaining low RRSE under moderate contamination and achieving lower errors than other methods around $r=20\%$, where its loss function mitigates the impact of moderate outliers. However, as contamination increases (e.g., $r>25\%$), its performance deteriorates. TARST, by comparison, maintains uniformly low RRSE values across nearly all conditions, showing minimal variation even under extreme settings (e.g., $r=50\%$, scale = 100). This indicates that the proposed thresholding strategy limits the effect of extreme observations by discarding outlier-dominated singular components, resulting in stable reconstruction performance across contamination conditions. These results suggest that TARST adapts well to situations where a small fraction of data exhibits highly dispersed values—analogous to heavy-tailed noise found in real-world datasets such as financial transactions, energy demand, or sensor readings. Unlike fixed-rank or iterative approaches, TARST automatically adjusts its rank to exclude such fluctuations without relying on an explicit noise model. This behavior implies that MP-based thresholding acts as a robustness regularizer, suppressing extreme singular components while preserving the global multilinear structure. Consequently, TARST demonstrates both noise resistance and resilience to heavy-tailed perturbations, bridging statistical robustness with computational tractability in tensor decomposition.
@@
Practically, TARST’s non-iterative and parameter-free design ensures reproducible results independent of analyst expertise or initialization. Its robustness to noise and outliers makes it suitable for various OR problems, including demand forecasting, anomaly detection, and efficiency evaluation involving multidimensional data. Furthermore, its simplicity and scalability enable deployment in large-scale, real-time analytical environments such as IoT and smart city monitoring.
@@
A current limitation of TARST is its assumption of i.i.d. Gaussian noise. Future research should extend the framework to handle non-Gaussian, anisotropic, or time-varying noise. Algorithmically, developing distributed or online implementations and incorporating adaptive mode-wise thresholding would improve scalability for streaming and large-scale data. From an application viewpoint, applying TARST to real-world datasets in domains such as transportation, finance, and energy, and integrating it with optimization-based decision-support systems represent promising directions. Additionally, establishing asymptotic guarantees for the estimated thresholds and comparing TARST’s statistical efficiency with other robust estimators would strengthen its theoretical foundation. In summary, TARST achieves theoretical optimality and computational efficiency while overcoming key limitations of conventional tensor decomposition methods, including rank specification, iterative optimization, and initialization sensitivity. The proposed framework provides a robust, reproducible, and theoretically grounded foundation for high-dimensional data analysis, advancing data-driven decision support in OR.
@@ 6 Conclusion
This study proposed TARST, a non-iterative and rank-free tensor decomposition method designed to overcome two major limitations of conventional approaches: the requirement for pre-specified ranks and the computational instability inherent in iterative optimization. By employing automatic thresholding based on the MP distribution, TARST achieves statistically valid rank estimation. Theoretical analysis using the WCTR further demonstrated that TARST guarantees computational tractability within polynomial time. Accordingly, TARST provides a unified methodological foundation for tensor decomposition that ensures both theoretical optimality and computational efficiency.
Simulation experiments demonstrated that TARST maintained the lowest reconstruction error across diverse noise intensities and outlier ratios, exhibiting remarkable stability compared with existing methods. These results indicate that the proposed approach is consistently robust to noise, outliers, and scale variation. In particular, its ability to estimate the underlying structure without rank specification or iterative refinement offers significant practical advantages.
From the perspective of OR, TARST contributes to both methodological innovation and practical decision support. Methodologically, it establishes a statistically grounded and computationally tractable framework for high-dimensional tensor analysis, bridging the gap between statistical estimation theory and algorithmic efficiency in OR. Practically, the robustness and parameter-free nature of TARST enable reliable, analyst-independent analysis in data-driven decision-support contexts such as demand forecasting, anomaly detection, and performance evaluation. Thus, TARST serves as a foundational analytical tool that connects theoretical rigor with operational applicability.
Future research directions include extending the framework to accommodate non-Gaussian and time-varying noise, developing distributed or online implementations, and integrating TARST with optimization algorithms to support real-time decision-making. Through these extensions, TARST has the potential to further bridge theory and practice in high-dimensional data analysis and contribute to the continued advancement of OR.